I have to make 50 cm cubed of sodium chloride that contains 0.2 moll How would i calculate how much sodium chloride is needed?
1
Expert's answer
2013-01-16T12:11:48-0500
You have 0.2 mol in 1000 cm^3 ( the same that 2 mol/L) and in 50 cm^3 you will have some other amount which is 0.2 * 50/1000 = 0.01 mol. You need to find mass of sodium chloride

m = n*Mw
Mw of NaCl is 58.44

m = 0.01*58.44 = 0.5844 g

So you need to mix 0.5844 g of NaCl and 50 cm^3 (50 ml) of water.